OpenStudy (anonymous):

There 6 entrances to the haunted house. If you do not exit the same way you entered, how many different ways can you enter and exit the house? do you just multiply 6 by the 5 remaing doors to get 30?

OpenStudy (anonymous):

6*5 = 35

OpenStudy (anonymous):

opps 6*5=30

OpenStudy (anonymous):

6*5 = 30

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

so am i correct?

OpenStudy (amriju):

of crse...u can enter n six ways..and each tym u getout u have 5 options...total 30

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

WELCOME